OpenStudy (peachtea):

The formula for the average of two numbers is m= a+b/2 Solve for a. a=______

KjSaif (kjsaif):

Flip the equation. a+1/2 b=m

KjSaif (kjsaif):

Add (-1)/2b to both sides.

KjSaif (kjsaif):

a + 1/2b + −1/2b = m + −1/2b

KjSaif (kjsaif):

solving that gets you the answer \[A = \frac{-1 }{ 2 }b+m\]

OpenStudy (peachtea):

Thank you so much for your help :D

KjSaif (kjsaif):

no problem!

KjSaif (kjsaif):

And i wanted to give you a big welcome so.........\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (peachtea):

Awe yay!! So colorful <3 Thanks so much!! Btw awesome profile pic hahahah c:

KjSaif (kjsaif):

lol thanks! :)

OpenStudy (peachtea):

Np :)